How do I solve this
It says solve by the substitution.
3y-x=0
x-4y=-2
my answer is 0 is this right? please help me.

Answer by suresh(20)

3y - x = 0 ------(1)
x - 4y = -2 -----(2)
From (1), we have x = 3y.------(3)
Plug this value in (2)
3y - 4y = -2
-y = -2
y = 2.
Plug y = 2 in (3), x = 3(2) = 6,
The solutions are x = 6 and y = 2.